To calculate the number average molar mass of a polymer sample with the following distribution of molar masses :
Nᵢ (Mᵢ/kg mol⁻¹)
5 1.000
10 2.000
10 50.000
10 10.000
5 20.000
1
Expert's answer
2020-12-24T03:53:02-0500

"M_n=\frac{\sum N_iM_i}{\sum N_i}=\frac{5*1+10*2+10*50+10*10+5*20}{5+10+10+10+5}=18.125" kg/mol